# Supplement

### v3

@@ 6 Discussion
We proceed by defining both of the penalty components $\mathbf{P}_{0},\mathbf{P}_{2}$ separately, as $\mathbf{P}_{\alpha}$ is just a linear combination of these matrices. Let $\mathbf{B}(t)=[b_{1}(t)|\ldots,|b_{Q}(t)]$ represent the basis functions, chosen to be orthonormal in $L^{2}([0,1])$. First, we define the zero-order penalty $\mathbf{P}_{0}$ element-wise. This derivation leverages the orthonormal definition of the $b_{i}(t)$.
[…]
=$=\begin{cases}1\text{ when }i=j\\ 0\text{ otherwise}\end{cases}$
The resulting matrix $\mathbf{P}_{0}=\mathbf{I}_{Q}$ by the definition of the basis $\mathbf{B}(t)$. This clearly indicates the role of this component (when non-zero) in ensuring the non-singularity of the final penalty $\mathbf{P}_{\alpha}$, similar to adding a ridge penalty in the context of regression.
[…]
We approximate these quantities using numerical integration based upon Gaussian quadrature.
[…]
S2 Joint and conditional posteriors
@@
$f(w_{\mu})\propto h_{\mu}^{\text{R}(\mathbf{P}_{\alpha})/2}\exp\left\{-\frac{h_{\mu}}{2}w^{t}_{\mu}\mathbf{P}_{\alpha}w_{\mu}\right\}$
@@
$\begin{split}&\prod_{i=1}^{N}MVN\{Y_{i}|\mathbf{B}(w_{\mu}+\sum_{k=1}^{K}\xi_{ik}\psi_{k}),\sigma^{2}I_{M}\}\times\Gamma^{-1}(\sigma^{2}|\alpha_{\sigma},\beta_{\sigma})\\ &\times\left\{\prod_{k=1}^{K}N(\xi_{ik}|0,\lambda_{k})\times\Gamma^{-1}(\lambda_{k}|\alpha_{\lambda},\beta_{\lambda})\times h_{k}^{\mathbf{R}(\mathbf{P}_{\alpha})/2}\exp\left(-\frac{h_{k}}{2}\psi_{k}^{t}\mathbf{P}_{\alpha}\psi_{k}\right)\times\Gamma(h_{k}|\alpha_{\psi},\beta_{\psi})\right\}\\ &\times h_{\mu}^{\mathbf{R}(\mathbf{P}_{\alpha})/2}\exp\left(-\frac{h_{\mu}}{2}w^{t}_{\mu}\mathbf{P}_{\alpha}w_{\mu}\right)\times\Gamma(h_{\mu}|\alpha_{\mu},\beta_{\mu})\times\mathbbm{1}(\lambda_{1}\geq\ldots\geq\lambda_{K})\times\mathbbm{1}(\mathbf{\Psi}\in\mathcal{V}_{Q,K})\end{split}$ (3)
@@
$[h_{\mu}|{\rm others}]\sim\Gamma(\mathbf{R}(\mathbf{P}_{\alpha})/2+\alpha_{\mu},w_{\mu}^{t}\mathbf{P}_{\alpha}w_{\mu}/2+\beta_{\mu})$
@@
$[h_{k}|{\rm others}]\sim\Gamma(\mathbf{R}(\mathbf{P}_{\alpha})/2+\alpha_{\psi},\psi_{k}^{t}\mathbf{P}_{\alpha}\psi_{k}/2+\beta_{\psi})$
@@
$\propto\prod_{i=1}^{N}(\sigma^{2})^{-M/2}\exp(-\frac{||Y_{i}-\mathbf{B}(w_{\mu}+\sum_{k=1}^{K}\xi_{ik}\psi_{k})||^{2}}{2\sigma^{2}})\times(\sigma^{2})^{-\alpha_{\sigma}-1}\exp(-\frac{\beta_{\sigma}}{\sigma^{2}})$
$\propto(\sigma^{2})^{-(NM/2+\alpha_{\sigma})-1}\exp\{-\frac{1}{\sigma^{2}}\left(\frac{1}{2}\sum_{i=1}^{n}||Y_{i}-\mathbf{B}(w_{\mu}+\sum_{k=1}^{K}\xi_{ik}\psi_{k})||^{2}+\beta_{\sigma}\right)\}$
@@
$[\sigma^{2}|{\rm others}]\sim\Gamma^{-1}(\frac{NM}{2}+\alpha_{\sigma},\frac{1}{2}\sum_{i=1}^{n}||Y_{i}-\mathbf{B}(w_{\mu}+\sum_{k=1}^{K}\xi_{ik}\psi_{k})||^{2}+\beta_{\sigma}\bigr{]})$
@@
The above joint distribution has the form of independent inverse gamma distributions ($[\lambda_{k}|{\rm others}]\sim\Gamma^{-1}(N/2+\alpha_{\lambda},\frac{1}{2}\sum_{i=1}^{N}\xi_{ik}^{2}+\beta_{\lambda})$), with the additional constraint that the $\lambda_{k}$ be ordered. Sampling from this type of joint distribution is possible using ordered transforms and corresponding Jacobian transforms as described in the STAN documentation (Team,, 2025).
@@
$\propto\prod_{i=1}^{N}\exp\left\{-\frac{||\mathbf{B}w_{\mu}-(Y_{i}-\mathbf{B}\sum_{k=1}^{K}\xi_{ik}\psi_{k})||^{2}}{2\sigma^{2}}-\frac{h_{\mu}}{2}w^{t}_{\mu}\mathbf{P}_{\alpha}w_{\mu}\right\}$
$\propto\exp\left\{-\sum_{i=1}^{N}\frac{||\mathbf{B}w_{\mu}||^{2}-2\langle Y_{i}-\mathbf{B}\sum_{k=1}^{K}\xi_{ik}\psi_{k},\mathbf{B}w_{\mu}\rangle}{2\sigma^{2}}-\frac{h_{\mu}}{2}w^{t}_{\mu}\mathbf{P}_{\alpha}w_{\mu}\right\}$
@@
$f(w_{\mu}|{\rm others})$ $\propto\exp\left\{-\sum_{i=1}^{N}\frac{||\mathbf{B}w_{\mu}||^{2}-2D_{i}^{t}\mathbf{B}w_{\mu}}{2\sigma^{2}}-w^{t}_{\mu}\frac{h_{\mu}\mathbf{P}_{\alpha}}{2}w_{\mu}\right\}$

### v4

@@ 6 Discussion
Considering the form of $\mathbf{P}_{\alpha}$, it becomes clear that $\mathbf{P}_{0}$ acts to ensure the non-singularity of the final penalty $\mathbf{P}_{\alpha}$ when $\alpha>0$, similar to adding a ridge penalty in the context of regression.
S2 Joint Prior on $h_{k}$ and $\boldsymbol{\Psi}$
We derive sufficient conditions under which the joint prior distribution on the smoothing parameters $h_{k}$ and the FPC spline coefficient matrix $\boldsymbol{\Psi}$ is proper. This is equivalent to showing under what assumptions it can be integrated to a constant, finite value. We assume here that ${\rm R}(\mathbf{P}_{\alpha})=Q$, as this is true for the tested implementation. Letting $K$ refer to the dimension of the FPC basis, $\Gamma(x|a,b)$ refer to a Gamma distribution with expectation $a/b$ evaluated at $x$, $\mathbf{H}={\rm diag}(h_{1},\ldots,h_{K})$, and ${\rm etr}(\cdot)$ refer to the exponential of the matrix trace, we proceed:
$g(\boldsymbol{\Psi},\mathbf{H})$ $=g_{\psi}(\boldsymbol{\Psi}|\mathbf{H})g_{h}(\mathbf{H})$
$\propto\frac{{\rm etr}(-\mathbf{H}\boldsymbol{\Psi}^{T}\mathbf{P}_{\alpha}\boldsymbol{\Psi}/2)}{{\rm Vol}(\mathcal{V}_{K,Q})}\times\prod_{i=1}^{K}h_{i}^{Q/2}\Gamma(h_{i}|\alpha_{\psi},\beta_{\psi})$
With the form of the density defined, we now integrate over the Stiefel manifold $\boldsymbol{\Psi}\in\mathcal{V}_{K,Q}$ and the Gamma-distributed smoothing parameters $h_{k}\geq 0$. Here, we use that $g_{\psi}(\boldsymbol{\Psi}|\mathbf{H})$ is proportional to a Matrix Bingham distribution with arguments $\mathbf{H}$ and $\mathbf{P}_{\alpha}/2$ (using commutativity of scalar and matrix multiplication). This distribution has normalizing constant $\Phi_{K,Q}(\mathbf{H},-\mathbf{P}_{\alpha}/2)^{-1}$, where $\Phi$ is the hypergeometric function of two matrix arguments. This function has the following series definition:
$\Phi_{K,Q}(\mathbf{H},-\mathbf{P}_{\alpha}/2)=\sum_{k=0}^{\infty}\frac{1}{k!}\sum_{|\kappa|=k}\frac{C_{\kappa}(\mathbf{H})C_{\kappa}(-\mathbf{P}_{\alpha}/2)}{C(I_{K})}$
where $C_{\kappa}()$ denotes the zonal polynomial related to partition $\kappa$ and the inner sum is over all partitions $\kappa$ of weight $k$ and length $\leq K$. Using this normalizing constant, we can integrate the joint prior density of the smoothing parameters and FPC spline coefficients as follows, letting $V={\rm Vol}(\mathcal{V}_{K,Q})$. Differentials are omitted for notational brevity.
$\int_{0}^{\infty}\cdots\int_{0}^{\infty}\int_{\boldsymbol{\Psi}\in\mathcal{V}_{K,Q}}g(\boldsymbol{\Psi},\mathbf{H})$ $=\int_{0}^{\infty}\cdots\int_{0}^{\infty}\int_{\boldsymbol{\Psi}\in\mathcal{V}_{K},Q}\frac{{\rm etr}(-\mathbf{H}\boldsymbol{\Psi}^{T}\mathbf{P}_{\alpha}\boldsymbol{\Psi}/2)}{V}\times\prod_{i=1}^{K}h_{i}^{Q/2}\Gamma(h_{i}|\alpha_{\psi},\beta_{\psi})$
$=\frac{1}{V}\int_{0}^{\infty}\cdots\int_{0}^{\infty}\Phi_{K,Q}(\mathbf{H},-\mathbf{P}_{\alpha}/2)\times\prod_{i=1}^{K}h_{i}^{Q/2}\Gamma(h_{i}|\alpha_{\psi},\beta_{\psi})$
$=\frac{1}{V}\int_{0}^{\infty}\cdots\int_{0}^{\infty}\sum_{k=0}^{\infty}\frac{1}{k!}\sum_{|\kappa|=k}\frac{C_{\kappa}(\mathbf{H})C_{\kappa}(-\mathbf{P}_{\alpha}/2)}{C_{\kappa}(I_{K})}\times\prod_{i=1}^{K}h_{i}^{Q/2}\Gamma(h_{i}|\alpha_{\psi},\beta_{\psi})$
At this point, we recall that the zonal polynomials are homogeneous (order equal to the weight $|\kappa|$), symmetric polynomials in the eigenvalues of the argument matrix. For positive definite matrix $\mathbf{P}_{\alpha}$, the eigenvalues of $\mathbf{P}_{\alpha}/2$ will be negative. The negative sign contribution from each such eigenvalue will be be present in every factor within each term of the polynomial, with cumulative power corresponding to the uniform order $|\kappa|$. This implies that $C_{\kappa}(-\mathbf{P}_{\alpha}/2)=(-1)^{|\kappa|}C_{\kappa}(\mathbf{P}_{\alpha}/2)$. Using this fact, the integral of interest becomes the following.
$\int_{0}^{\infty}\cdots\int_{0}^{\infty}\int_{\mathcal{V}_{K,Q}}g(\boldsymbol{\Psi},\mathbf{H})$ $=\frac{1}{V}\int_{0}^{\infty}\cdots\int_{0}^{\infty}\sum_{k=0}^{\infty}\frac{(-1)^{k}}{k!}\sum_{|\kappa|=k}\frac{C_{\kappa}(\mathbf{H})C_{\kappa}(\mathbf{P}_{\alpha}/2)}{C_{\kappa}(I_{K})}\times\prod_{i=1}^{K}h_{i}^{Q/2}\Gamma(h_{i}|\alpha_{\psi},\beta_{\psi})$
We now recall the combinatorial definition of zonal polynomials. In particular, each zonal polynomial $C_{\kappa}(A)$ for partition $\kappa$ and matrix $A\in m\times m$ can be represented as a linear combination of the monomial symmetric functions $M_{\lambda}(\cdot)$ in the eigenvalues of AA $(\sigma_{1}\ldots,\sigma_{m})$:
$C_{\kappa}(A)=\sum_{\lambda\leq\kappa}c_{\kappa,\lambda}M_{\lambda}(\sigma_{1},\ldots,\sigma_{m})$
These monomial symmetric functions $M_{\lambda}(\cdot)$ correspond to partitions $\lambda$ which have equal weight to $(|\lambda|=|\kappa|)$ and are before $\kappa$ according to lexicographic ordering. All coefficients $c_{\kappa,\lambda}$ are non-negative. The monomial symmetric functions are defined as follows for partition $\lambda$ of length $\mathcal{L}$.
$M_{\lambda}(\sigma_{1},\ldots,\sigma_{m})=\sum_{p\in P}\sigma_{p_{1}}^{\lambda_{1}}\cdots\sigma_{p_{\mathcal{L}}}^{\lambda_{\mathcal{L}}}$
where the summation is taken over all distinct permutations $p\in P$ of $\mathcal{L}$ elements drawn from $\{1,\ldots,m\}$ (Bagyan and Richards,, 2024). By this definition, it follows that the zonal polynomial of a positive semi-definite matrix will be positive when the length of the partition is at most the rank of the matrix, taking zero-value when the length is greater. Note that the summation over partitions $\kappa$ above requires that all have length $\leq K<Q$.
As $\mathbf{P}_{\alpha}/2$ and $I_{K}$ are positive definite, it follows that $C_{\kappa}(\mathbf{P}_{\alpha}/2)>0$ and $C_{\kappa}(I_{K})>0$. Similarly, $\mathbf{H}$ being positive semi-definite implies that $C_{\kappa}(\mathbf{H})\geq 0$. This, combined with the non-negativity of each factor $\frac{h_{i}^{Q/2}}{k!}\Gamma(h_{i}|\alpha_{\psi},\beta_{\psi})$, implies that the sign of the series indexed by $k$ is entirely decided by the $(-1)^{k}$ term. Using this fact, we can break the summation into positive and negative portions.
$\int_{0}^{\infty}\cdots\int_{0}^{\infty}\int_{\mathcal{V}_{K,Q}}g(\boldsymbol{\Psi},\mathbf{H})$ $=\frac{1}{V}\int_{0}^{\infty}\cdots\int_{0}^{\infty}\sum_{k=0}^{\infty}\frac{1}{(2k)!}\sum_{|\kappa|=2k}\frac{C_{\kappa}(\mathbf{H})C_{\kappa}(\mathbf{P}_{\alpha}/2)}{C_{\kappa}(I_{K})}\times\prod_{i=1}^{K}h_{i}^{Q/2}\Gamma(h_{i}|\alpha_{\psi},\beta_{\psi})$
$-\frac{1}{V}\int_{0}^{\infty}\cdots\int_{0}^{\infty}\sum_{k=0}^{\infty}\frac{1}{(2k+1)!}\sum_{|\kappa|=2k+1}\frac{C_{\kappa}(\mathbf{H})C_{\kappa}(\mathbf{P}_{\alpha}/2)}{C_{\kappa}(I_{K})}\times\prod_{i=1}^{K}h_{i}^{Q/2}\Gamma(h_{i}|\alpha_{\psi},\beta_{\psi})$
We will now apply Tonelli’s theorem to each part, once for each Gamma-distributed smoothing parameter. To verify that Tonelli’s theorem applies, we consider the necessary assumptions: (1) the measure spaces integrated over must be $\sigma$-finite, (2) the integrand must be non-negative and, (3) the integrand must be measurable. Each application of Tonelli’s theorem here applies to a function $f(h_{j},k)$ as defined for the first and second terms below. We use here that zonal polynomials are polynomial functions of the eigenvalues, which are just the diagonal elements for a diagonal matrix such as $\mathbf{H}$.
$\text{Term 1: }f(h_{j},k)$ $=\frac{1}{(2k)!}\sum_{|\kappa|=2k}\frac{C_{\kappa}(h_{1},\ldots,h_{j},\ldots,h_{K})C_{\kappa}(\mathbf{P}_{\alpha}/2)}{C_{\kappa}(I_{K})}\times\prod_{i=1}^{K}h_{i}^{Q/2}\Gamma(h_{i}|\alpha_{\psi},\beta_{\psi})$
$\text{Term 2: }f(h_{j},k)$ $=\frac{1}{(2k+1)!}\sum_{|\kappa|=2k+1}\frac{C_{\kappa}(h_{1},\ldots,h_{j},\ldots,h_{K})C_{\kappa}(\mathbf{P}_{\alpha}/2)}{C_{\kappa}(I_{K})}\times\prod_{i=1}^{K}h_{i}^{Q/2}\Gamma(h_{i}|\alpha_{\psi},\beta_{\psi})$
We consider the $f(h_{j},k)$ for both terms concurrently. In these functions, $h_{j}\in(\mathbb{R}^{+},\mathbf{B}(\mathbb{R}^{+}),\gamma)$ for probability measure $\gamma$ corresponding to the Gamma distribution and $k\in(\mathbb{N},2^{\mathbb{N}},\mu)$ for counting measure $\mu$. To assumption (1), probability measures and the counting measure on $\mathbb{N}$ are both known to be $\sigma$-finite by definition. To assumption (2), we have already demonstrated that $f(h_{j},k)$ in both cases is non-negative using the properties of zonal polynomials for positive semi-definite matrices. For the final assumption (3), note that it suffices to show that the $f(h_{j},k)$ are Borel-measurable for fixed $k$ (see Bartle, (1995) chapter 10). This follows directly from the fact that each term will reduce to an exponential polynomial in $h_{j}$, which are continuous and thus measurable. With this in mind, we proceed:
$\int_{0}^{\infty}\cdots\int_{0}^{\infty}\int_{\mathcal{V}_{K,Q}}g(\boldsymbol{\Psi},\mathbf{H})$ $=\frac{1}{V}\sum_{k=0}^{\infty}\frac{1}{(2k)!}\sum_{|\kappa|=2k}\frac{C_{\kappa}(\mathbf{P}_{\alpha}/2)}{C_{\kappa}(I_{K})}\times\int_{0}^{\infty}\cdots\int_{0}^{\infty}C_{\kappa}(\mathbf{H})\prod_{i=1}^{K}h_{i}^{Q/2}\Gamma(h_{i}|\alpha_{\psi},\beta_{\psi})$
$-\frac{1}{V}\sum_{k=0}^{\infty}\frac{1}{(2k+1)!}\sum_{|\kappa|=2k+1}\frac{C_{\kappa}(\mathbf{P}_{\alpha}/2)}{C_{\kappa}(I_{K})}\times\int_{0}^{\infty}\cdots\int_{0}^{\infty}C_{\kappa}(\mathbf{H})\prod_{i=1}^{K}h_{i}^{Q/2}\Gamma(h_{i}|\alpha_{\psi},\beta_{\psi})$
At this point, we are able to combine the difference back into a single sum, resulting in the following:
$\int_{0}^{\infty}\cdots\int_{0}^{\infty}\int_{\mathcal{V}_{K,Q}}g(\boldsymbol{\Psi},\mathbf{H})$ $=\frac{1}{V}\sum_{k=0}^{\infty}\frac{(-1)^{k}}{k!}\sum_{|\kappa|=k}\frac{C_{\kappa}(\mathbf{P}_{\alpha}/2)}{C_{\kappa}(I_{K})}\times\int_{0}^{\infty}\cdots\int_{0}^{\infty}C_{\kappa}(\mathbf{H})\prod_{i=1}^{K}h_{i}^{Q/2}\Gamma(h_{i}|\alpha_{\psi},\beta_{\psi})$
To prove that this alternating series converges, it is sufficient to prove absolute convergence. We define the series of the absolute values below using $S_{A}(\alpha_{\psi},\beta_{\psi})$ – a function of the smoothing prior parameters $\alpha_{\psi},\beta_{\psi}$. If this series converges, then the integral $\int_{0}^{\infty}\cdots\int_{0}^{\infty}\int_{\mathcal{V}_{K,Q}}g(\boldsymbol{\Psi},\mathbf{H})$ will be finite as desired.
$S_{A}(\alpha_{\psi},\beta_{\psi})$ $=\frac{1}{V}\sum_{k=0}^{\infty}\frac{1}{k!}\sum_{|\kappa|=k}\frac{C_{\kappa}(\mathbf{P}_{\alpha}/2)}{C_{\kappa}(I_{K})}\times\int_{0}^{\infty}\cdots\int_{0}^{\infty}C_{\kappa}(\mathbf{H})\prod_{i=1}^{K}h_{i}^{Q/2}\Gamma(h_{i}|\alpha_{\psi},\beta_{\psi})$
We can now apply inequality $5.26$ from Bagyan and Richards, (2024), derived from Faraut and Korányi, (1994). This step will provide an upper bounding series for the non-negative $S_{A}(\alpha_{\psi},\beta_{\psi})$ which is easier to work with. This inequality states that, for $\Sigma$ being a $d\times d$ positive definite matrix with ordered eigenvalues $\sigma_{1}\geq\ldots\geq\sigma_{d}$, it follows that
$C_{\kappa}(\Sigma)\leq C_{\kappa}(I_{d})\prod_{j=1}^{d}\sigma_{j}^{\kappa_{j}}$
where the partition $\kappa$ is permitted to contain zero entries. The diagonal matrix $\mathbf{H}$ will have eigenvalues equal to its diagonal elements $h_{k}$, which will be non-zero with probability 1, meaning $\mathbf{H}$ will be positive definite with probability 1. This almost-sure inequality of the integrands will produce the proper inequality when taking the requisite expectation/integral.
$S_{A}(\alpha_{\psi},\beta_{\psi})$ $\leq\frac{1}{V}\sum_{k=0}^{\infty}\frac{1}{k!}\sum_{|\kappa|=k}\frac{C_{\kappa}(I_{K})C_{\kappa}(\mathbf{P}_{\alpha}/2)}{C_{\kappa}(I_{K})}\times\int_{0}^{\infty}\cdots\int_{0}^{\infty}\prod_{i=1}^{K}h_{i}^{\kappa_{i}+Q/2}\Gamma(h_{i}|\alpha_{\psi},\beta_{\psi})$
$=\frac{1}{V}\sum_{k=0}^{\infty}\frac{1}{k!}\sum_{|\kappa|=k}C_{\kappa}(\mathbf{P}_{\alpha}/2)\prod_{i=1}^{K}\mathbb{E}[h_{i}^{\kappa_{i}+Q/2}]$
We have reduced the integral of interest to the independent product of moments of each smoothing coefficient. Applying the known form of these moments in terms of the shape $\alpha_{\psi}$ and rate $\beta_{\psi}$ results in the following:
$S_{A}(\alpha_{\psi},\beta_{\psi})$ $\leq\frac{1}{V}\sum_{k=0}^{\infty}\frac{1}{k!}\sum_{|\kappa|=k}C_{\kappa}(\mathbf{P}_{\alpha}/2)\prod_{i=1}^{K}\frac{\Gamma(\alpha_{\psi}+\kappa_{i}+Q/2)}{\beta_{\psi}^{\kappa_{i}+Q/2}\Gamma(\alpha_{\psi})}$
$=\frac{1}{V}\sum_{k=0}^{\infty}\frac{1}{k!\times\beta_{\psi}^{k+KQ/2}}\sum_{|\kappa|=k}C_{\kappa}(\mathbf{P}_{\alpha}/2)\prod_{i=1}^{K}\prod_{j=1}^{\kappa_{i}+Q/2}(\alpha_{\psi}+j-1)$
We will now consider the last factor in the above series, for which we provide the shorthand $Pr(\kappa)=\prod_{i=1}^{K}\prod_{j=1}^{\kappa_{i}+Q/2}(\alpha_{\psi}+j-1)$. We endeavor to demonstrate that $Pr(\kappa)$ is maximized by the partition which places all of the weight on a single entry, denoted $\kappa=(k)$ in literature. Towards this end, we will demonstrate that $Pr(\kappa)$ for $\kappa\neq(k)$ is always strictly increased by moving weight to the first entry from the last non-zero entry. Before proceeding, it is important to recall that each $\kappa$ is defined such that the entries are non-decreasing: $\kappa_{1}\geq\kappa_{2}\geq\ldots\geq\kappa_{K}\geq 0$.
Consider an arbitrary integer partition $\kappa\neq(k)$ with length (number of non-zero components) $\mathcal{L}(\kappa)$. We compare $Pr(\kappa)$ with $Pr(\kappa^{\prime})$ for $\kappa^{\prime}_{1}=\kappa_{1}+1,\kappa^{\prime}_{i}=\kappa_{i}$ when $1<i<\mathcal{L}(\kappa)$, and $\kappa^{\prime}_{\mathcal{L}(\kappa)}=\kappa_{\mathcal{L}(\mathcal{\kappa})}-1$. This $\kappa^{\prime}$ still has the same weight, and should satisfy the ordering requirement, so it is a well-defined partition. We can now calculate $\frac{Pr(\kappa^{\prime})}{Pr(\kappa)}$ as follows. This ratio will always be well-defined, as $Pr(\cdot)$ will always be a product of positive factors.
$\frac{Pr(\kappa^{\prime})}{Pr(\kappa)}$ $=\frac{\prod_{i=1}^{K}\prod_{j=1}^{\kappa^{\prime}_{i}+Q/2}(\alpha_{\psi}+j-1)}{\prod_{i=1}^{K}\prod_{j=1}^{\kappa_{i}+Q/2}(\alpha_{\psi}+j-1)}$
$=\frac{\prod_{j=1}^{(\kappa_{1}+1)+Q/2}(\alpha_{\psi}+j-1)}{\prod_{j=1}^{\kappa_{1}+Q/2}(\alpha_{\psi}+j-1)}\times\frac{\prod_{j=1}^{(\kappa_{\mathcal{L}(\kappa)}-1)+Q/2}(\alpha_{\psi}+j-1)}{\prod_{j=1}^{\kappa_{\mathcal{L}(\kappa)}+Q/2}(\alpha_{\psi}+j-1)}$
$=\frac{\alpha_{\psi}+\kappa_{1}+Q/2}{\alpha_{\psi}+\kappa_{\mathcal{L}(\kappa)}+Q/2-1}$
As $\kappa_{1}\geq\kappa_{\mathcal{L(\kappa)}}$ inherently, it follows that the above ratio will be >1>1, so $Pr(\cdot)$ is strictly increased by moving weight to the first entry. Then, beginning at any arbitrary partition, transitioning weight from the last non-zero entry in the partition to the first will always increase $Pr(\cdot)$. As any partition for can be related to $(k)$ through a series of such ”transitions”, it follows from the transitive property that $Pr[(k)]>Pr(\kappa)$ for arbitrary partition $\kappa$.
Using the above result, we establish another bounding Series on $S_{A}(\alpha_{\psi},\beta_{\psi})$. Through this step, we are able to move the $Pr(\cdot)$ term out of the summation over partitions of weight $k$, allowing use of the identity $\sum_{|\kappa|=k}C_{\kappa}(A)={\rm tr}(A)^{k}$ (Bagyan and Richards,, 2024).
$S_{A}(\alpha_{\psi},\beta_{\psi})$ $\leq\frac{1}{V}\sum_{k=0}^{\infty}\frac{1}{k!\times\beta_{\psi}^{k+KQ/2}}\sum_{|\kappa|=k}C_{\kappa}(\mathbf{P}_{\alpha}/2)Pr(\kappa)$
$<\frac{1}{V}\sum_{k=0}^{\infty}\frac{1}{k!\times\beta_{\psi}^{k+KQ/2}}\sum_{|\kappa|=k}C_{\kappa}(\mathbf{P}_{\alpha}/2)Pr[(k)]$
$=\frac{1}{V}\sum_{k=0}^{\infty}\frac{1}{k!\times\beta_{\psi}^{k+KQ/2}}\left[\sum_{|\kappa|=k}C_{\kappa}(\mathbf{P}_{\alpha}/2)\right]Pr[(k)]$
$=\frac{1}{V}\sum_{k=0}^{\infty}\frac{1}{k!\times\beta_{\psi}^{k+KQ/2}}{\rm tr}(\mathbf{P}_{\alpha}/2)^{k}Pr[(k)]$
$=\frac{1}{V}\sum_{k=0}^{\infty}\frac{1}{k!\times\beta_{\psi}^{k+KQ/2}\times 2^{k}}{\rm tr}(\mathbf{P}_{\alpha})^{k}Pr[(k)]$
We can now discern under what conditions this bounding series, and thus the original series by comparison test, converge. This is accomplished below using the ratio test.
$a_{k}$ $=\frac{1}{k!\times\beta_{\psi}^{k+KQ/2}\times 2^{k}}{\rm tr}(\mathbf{P}_{\alpha})^{k}Pr[(k)]$
$\implies\frac{a_{k+1}}{a_{k}}$ $=\frac{\frac{1}{(k+1)!\times\beta_{\psi}^{(k+1)+KQ/2}\times 2^{k+1}}{\rm tr}(\mathbf{P}_{\alpha})^{(k+1)}Pr[(k+1)]}{\frac{1}{k!\times\beta_{\psi}^{k+KQ/2}\times 2^{k}}{\rm tr}(\mathbf{P}_{\alpha})^{k}Pr[(k)]}$
$=\frac{{\rm tr}(\mathbf{P}_{\alpha})}{2(k+1)\beta_{\psi}}\times\frac{\prod_{j=1}^{(k+1)+Q/2}(\alpha_{\psi}+j-1)\times\prod_{i=2}^{K}\prod_{p=1}^{Q/2}(\alpha_{\psi}+p-1)}{\prod_{j^{\prime}=1}^{k+Q/2}(\alpha_{\psi}+j^{\prime}-1)\times\prod_{i^{\prime}=2}^{K}\prod_{p^{\prime}=1}^{Q/2}(\alpha_{\psi}+p^{\prime}-1)}$
$=\frac{{\rm tr}(\mathbf{P}_{\alpha})(\alpha_{\psi}+k+Q/2)}{2(k+1)\beta_{\psi}}$
Taking the limit as $k\rightarrow\infty$ provides the following.
$\lim_{k\rightarrow\infty}\frac{a_{k+1}}{a_{k}}$ $=\lim_{k\rightarrow\infty}\frac{{\rm tr}(\mathbf{P}_{\alpha})(\alpha_{\psi}+k+Q/2)}{2(k+1)\beta_{\psi}}$
$=\frac{{\rm tr}(\mathbf{P}_{\alpha})}{2\beta_{\psi}}\lim_{k\rightarrow\infty}\frac{\alpha_{\psi}+k+Q/2}{k+1}$
$=\frac{{\rm tr}(\mathbf{P}_{\alpha})}{2\beta_{\psi}}$
By the ratio test, the above series $a_{k}$ converges if and only if $\lim_{k\rightarrow\infty}\frac{a_{k+1}}{a_{k}}<1$. This occurs precisely when $\frac{{\rm tr}(\mathbf{P}_{\alpha})}{2\beta_{\psi}}<1\implies{\rm tr}(\mathbf{P}_{\alpha})/2<\beta_{\psi}$. This informs our choice of the prior parameter $\beta_{\psi}$, as it must be greater than the half of the trace (or half of the sum of the eigenvalues) for the penalty matrix. This makes some intuitive sense, as the smoothing parameters having overly heavy upper tails, combined with there being non-trivial regions of the manifold $\mathcal{V}_{K,Q}$ over which penalization is quite extreme (quantified through the eigenvalues themselves), could result in an unbounded integrated penalty.
S3 Joint and conditional posteriors
@@
$f(w_{\mu})\propto h_{\mu}^{\mathbf{R}(\mathbf{P}_{\alpha})/2}\exp\left\{-\frac{h_{\mu}}{2}w^{t}_{\mu}\mathbf{P}_{\alpha}w_{\mu}\right\}$
@@
$\begin{split}&\prod_{i=1}^{N}MVN\{Y_{i}|\mathbf{B}(w_{\mu}+\sum_{k=1}^{K}\xi_{ik}\psi_{k}),\sigma^{2}I_{M}\}\times\Gamma^{-1}(\sigma^{2}|\alpha_{\sigma},\beta_{\sigma})\\ &\times\left\{\prod_{k=1}^{K}N(\xi_{ik}|0,\lambda_{k})\times\Gamma^{-1}(\lambda_{k}|\alpha_{\lambda},\beta_{\lambda})\times h_{k}^{\mathbf{R}(\mathbf{P}_{\alpha})/2}\exp\left(-\frac{h_{k}}{2}\psi_{k}^{t}\mathbf{P}_{\alpha}\psi_{k}\right)\times\Gamma(h_{k}|\alpha_{\psi},\beta_{\psi})\right\}\\ &\times h_{\mu}^{\mathbf{R}(\mathbf{P}_{\alpha})/2}\exp\left(-\frac{h_{\mu}}{2}w^{t}_{\mu}\mathbf{P}_{\alpha}w_{\mu}\right)\times\Gamma(h_{\mu}|\alpha_{\mu},\beta_{\mu})\times\mathbbm{1}(\lambda_{1}\geq\ldots\geq\lambda_{K})\times\mathbbm{1}(\mathbf{\Psi}\in\mathcal{V}_{Q,K})\end{split}$
@@
$h_{\mu}|{\rm others}]\sim\Gamma(\mathbf{R}(\mathbf{P}_{\alpha})/2+\alpha_{\mu},w_{\mu}^{t}\mathbf{P}_{\alpha}w_{\mu}/2+\beta_{\mu})$
@@
$h_{k}|{\rm others}]\sim\Gamma(\mathbf{R}(\mathbf{P}_{\alpha})/2+\alpha_{\psi},\psi_{k}^{t}\mathbf{P}_{\alpha}\psi_{k}/2+\beta_{\psi})$
@@
[…]
@@
$\sigma^{2}|{\rm others}]\sim\Gamma^{-1}(\frac{NM}{2}+\alpha_{\sigma},\frac{1}{2}\sum_{i=1}^{n}||Y_{i}-\mathbf{B}(w_{\mu}+\sum_{k=1}^{K}\xi_{ik}\psi_{k})||^{2}+\beta_{\sigma}\bigr])$
@@
The above joint distribution has the form of independent inverse gamma distributions ($\lambda_{k}|{\rm others}]\sim\Gamma^{-1}(N/2+\alpha_{\lambda},\frac{1}{2}\sum_{i=1}^{N}\xi_{ik}^{2}+\beta_{\lambda})$), with the additional constraint that the $\lambda_{k}$ be ordered. Sampling from this type of joint distribution is possible using ordered transforms and corresponding Jacobian transforms as described in the STAN documentation (Team,, 2025).
@@
[…]
@@
[…]